OpenStudy (anonymous):

The length of a rectangle is 10 more than twice its width. The area is 72square inches. Find the dimensions of the rectangle

OpenStudy (anonymous):

area= length x width let x be the width area= (10x)(x) \(72=10x^2\) and solve for 'x' that will be the width substitute the value of 'x' in length=10x

OpenStudy (anonymous):

Thank you Data_lg2

OpenStudy (anonymous):

But can you explain why it's not 2w-10=72^2in

OpenStudy (anonymous):

np :) and \(\Large \color{blue}{Welcome\ to\ OPENSTUDY!! ☺☺☺}\) hmm.. my bad, i did not read the question properly "length of a rectangle is 10 more than twice its width" so lemme do it again let x-width \(\Large 10+2x = 72\) ^this should be the equation, sorry:/ you almost got it, but instead of "-", it should be "+" because it says there that it is "10 more than" (^_^)

OpenStudy (anonymous):

Thank you again And fixing your mistake

OpenStudy (anonymous):

hehehe, sorry :/